andy would like to put a row of decorative tile around his circular pole. the radius of a pole is 5 feet. if the tile cost 12 dollars per foot how much will the tile cost?
Answer by checkley77(12569)

C=2pi*r
C=2*3.14*5
C=2*15.7
C=31.4 ft.
31.4*12=$376.8 cost of the tile.